OpenStudy (anonymous):

a=2 (b+c) solve for c

OpenStudy (anonymous):

(a/2) -c

OpenStudy (anonymous):

er, -b. My bad. It should be (a/2) - b = c

OpenStudy (anonymous):

a/2 - b

OpenStudy (anonymous):

a = 2b + 2c transpose 2b to left hand side a-2b = 2c divide both sides by 2 (a-2b)/2 = 2c/c (a-2b)/2 = c

OpenStudy (anonymous):

Kushashwa, but then you could just break it apart, and simplify it

OpenStudy (anonymous):

yes yes u r rigt mknight but the answer i found is also correct but u r also right

OpenStudy (anonymous):

yay! correctness!